X/250The local phone company is offering a phone line plan that costs a fixed monthly charge plus a long distance rate per minute. The plan charges $83.00 for 100 minutes of a long distance calling. For $95.00, you get 500 minutes of long distance calling. Write a linear model for the cost c(in dollars) of the plan for 250 minutes of long distance calling.
Answer by checkley77(12844)

(100,83) & (500,95)
RATE LINE=(95-83)/(500-100)
RL=12/400
RL=.03
Y=mX+b
83=.03*100+b
83=3+b
b=83-3
b=80
Y=.03X+b
SETTING Y=250 WE SOLVE FOR X.
Y=.03*250+80
Y=7.5+80
Y=7.5+80
Y=$87.5 FOR 250 MINUTES.